# Machine-Learned Potentials for Solvation Modeling

### v2

@@ III.9   Error metrics
[…]
Beyond direct error metrics, models are evaluated based on their ability to reproduce experimentally observable or trajectory-derived quantities, such as solvation free energies ($\Delta G_{\text{solv}}$), RDFs, hydration numbers, or diffusivities. Both energy and force accuracies contribute to the reliability of these higher-level properties, with force fidelity especially important for trajectory-based observables.
@@ III.10   Deployment in solvation modeling
A variety of open-source tools support the training and deployment of MLPs in solvation modeling, some of which are collected in Table 6. The recent editorial survey Rupp, Küçükbenli, and Csányi (2024) highlights community-wide efforts to develop software ecosystems for atomistic ML, with a specific focus on the integration of MLPs into the current tooling landscape, which includes simulation engines, workflow automation, and dataset infrastructure.
@@
Reproducibility in MLP-aided solvation modeling depends on detailed reporting of training data, model architectures, simulation parameters, and validation metrics. Beyond ML benchmarks, rigorous quantum chemical benchmarking also plays a key role in validating interaction energies in solvated systems. Chen et al. Chen et al. (2020) systematically assessed DFT, composite, and DLPNO-CCSD(T)/CBS methods on glycine–water clusters up to 40 waters, showing that range-separated hybrids like B3LYP-D3(BJ) and $\omega$B97M-V achieve MADs of 2–4 kJ/mol relative to high-level wavefunction theory. These error margins serve as a reference for validating MLP predictions on microsolvated clusters. Transfer learning from quantum chemistry to experiment has also emerged as a strategy for mitigating data scarcity and experimental noise in solvation modeling. Vermeire and Green Vermeire and Green (2021) demonstrated that pretraining MP-NN-MLPs on COSMO-RS–derived solvation free energies (CombiSolv-QM) and fine-tuning on curated experimental data (CombiSolv-Exp) can reduce test MAEs from  0.47 to  0.21 kcal/mol, and as low as 0.09 kcal/mol on high-accuracy subsets.
@@ IV   Case Studies: Trends and Themes
    A2: Direct prediction of solvation-dependent properties without PES construction. In these studies, MLPs are not used to model full PESs. Instead, they are trained to predict solvation-dependent properties—such as dipole moments, redox potentials, or solvation energies—directly from the molecular structure or other structural descriptors. These approaches bypass full force field construction and focus on property-specific inference.
@@ IV.1.1   Water contact layer at solid/liquid interfaces modeled by NN-MLPs
Gäding et al. Gäding et al. (2024) applied MD simulations powered by committee neural network potentials (C-NNPs) to investigate the structure of interfacial water at six different solid surfaces: graphene, MoS2, Au(111), Au(100), Pt(111), and Pt(100). The models were built on the BPNN architecture, using ACSFs as descriptors, and were trained iteratively on reference ab initio data using an active learning protocol enhanced by a query-by-committee (QbC) approach. To monitor model reliability during training, the authors evaluated the RMSE in the potential energy per atom.
@@ IV.1.3   Metadynamics of oxygen reduction at Au–water interfaces
To ensure broad coverage of chemical space, CUR matrix decomposition—a low-rank approximation technique to decompose a matrix $\mathbf{A}$ into a compressed matrix product $\mathbf{C}\times\mathbf{U}\times\mathbf{R}$—was employed during dataset curation, with Euclidean distances in structural space used to identify and retain the most diverse and informative samples. The resulting model achieved energy MAEs below 1 meV/atom and force MAEs between 12.6 and 16.3 meV/Å.
@@ IV.1.4   Neural network potential for protonated water clusters in helium droplets
Schran et al. Schran et al. (2018) trained an BPNN-MLP model to simulate helium–solute interactions in protonated water clusters (H3O+ and H5O+${}_{2}^{+}$), spanning both microsolvated and bulk regimes. The NN-MLP was initially trained on two-body interaction energies computed at the coupled cluster singles doubles with perturbative triples (CCSD(T)) level using counterpoise-corrected aug-cc-pVTZ basis sets, ensuring high fidelity to electronic structure data. Radial and angular symmetry functions were employed to preserve translational, rotational, and permutational invariances. When validated against CCSD(T) energy grids for fixed solute geometries, the NN-MLP achieved a small mean absolute deviation (MAD) of about 0.04 kJ/mol.
To incorporate nuclear quantum effects (NQEs), the trained NNP was coupled to ab initio path integral molecular dynamics (AI-PIMD) and path integral Monte Carlo (PIMC) simulations at 1.67 K. The AI-PIMD simulations employed the PBE exchange–correlation functional with D3 dispersion corrections for up to two-body terms. Simulations with the NN-MLP accurately captured structural observables such as spatial distribution functions (SDFs) and RDFs of helium atoms surrounding the solutes, both for microsolvated states and extended liquid helium environments.
To improve coverage of the PES, a refinement step was introduced by analyzing the symmetry function values from NN-MLP-driven path integral molecular dynamics (PIMD) trajectories and identifying structural regions underrepresented in the original training set. For H3O+ and H5O+${}_{2}^{+}$, 3545 and 4101 such structures were retrieved, respectively. These were recomputed at the CCSD(T) level and added to the training dataset, further refining the potential and enhancing its accuracy.
@@
Interestingly, for the Zundel cation (H5O+${}_{2}^{+}$) with asymmetric proton sharing, the SDFs showed nontrivial anomalies in helium placement. The helium atoms were displaced due to competing minima on the NN-MLP grid, leading to a slightly shallower interaction potential near the excess proton. Nonetheless, the model captured key structural motifs and reproduced solvation features across a wide range of system sizes, from isolated complexes to bulk-like conditions, demonstrating its versatility and transferability.
@@ IV.1.5   Solvent-driven ring opening of N-enoxyphthalimide
Célerse et al. Célerse et al. (2024) developed a DeepMD MLP using the DeepPot-SE framework and data at the hybrid-DFT-level with transfer learning to investigate solvent-mediated reactivity in bulk-phase environments. Their workflow combined active learning with well-tempered metadynamics (WT-MTD) to construct fully converged free energy surface (FES) s for the ring-opening reaction of N-enoxyphthalimide in methanol and trifluoroethanol (TFE). These simulations uncovered a diverse ensemble of transition state (TS) configurations, often stabilized by up to five explicit solvent molecules, reflecting the inherently dynamic nature of solute–solvent interactions.
@@ IV.1.6   Explicit solvent modeling of adsorption and reactivity at metal–water interfaces
Chen et al. Chen, Zhang, and Zhang (2023) demonstrated the power of machine mearning-accelerated molecular dynamics (MLaMD) with active learning to model heterogeneous catalysis under explicit solvation conditions. They employed an MTP-based linear-MLP, trained via Maxvol-based active learning. The resulting model retained DFT-level accuracy (MADs of  1 meV/atom and 0.04 eV/Å). MTP outperformed SchNet and ANI in predicting water energetics and RDFs, achieving better accuracies even with small training sets ( 800 DFT configurations) and enabling long-timescale MLaMD simulations of up to 500 ps compared to the AIMD simulations (50 ps).
@@ IV.1.8   Explicit solvent dynamics of the Menshutkin reaction via DeepPot-SE
WT-MetaD simulations using the trained NN-MLP allowed reconstruction of converged FES, yielding a barrier of 29.7 kcal/mol and a reaction free energy of $-39\pm 2$ kcal/mol, consistent with experimental trends and outperforming SMD-based implicit solvation. FES analysis revealed significant changes in solvation structure along the reaction coordinate, particularly enhanced stabilization of the charged products via hydrogen bonding. The study highlights how MLPs trained with explicit water environments can capture both the energetic and dynamic solvation effects that are missing in continuum models.
@@
Spatial distribution functions revealed two prominent solvation caps where $\pi$-hydrogen bonds form, with comparable structures in both solvents. A time-resolved existence criterion was used to determine bond populations and lifetimes (1.7–1.8 ps), and vibrational spectroscopy (via TRPMD) captured blue-shifted OH stretching features in water, consistent with experimental Raman SC spectra. Importantly, water showed significant $\pi$-hydrogen bond anticooperativity between the two binding sites on benzene, while ammonia did not, due to weaker solvent–solvent H-bonding. Despite accurate reproduction of most properties, the C-NNP model underestimated the anticooperativity effect, highlighting limitations of local descriptors and standard training protocols. The authors suggest using MP-NN-MLPs architectures such as NequIP with customized active learning to address such long-range effects.
@@ IV.1.10   Water at solid–liquid and confined interfaces
The C-NNPs achieved high accuracy with compact training sets ($\sim$300 configurations) and were validated using an automated protocol analyzing RDFs, VDOS, and force RMSEs. For the water–TiO2 interface, long (5 ns) simulations enabled insight into interfacial layering, diffusion anisotropy, and free energy profiles, revealing strongly structured and immobile water contact layers. The framework leverages symmetry functions and fixed hyperparameters, requiring minimal user tuning, and is implemented in the open-source packages. This study exemplifies a scalable workflow for solvation-aware MLP development and application.
@@ IV.2.2   ML-aided computation of absolute redox potentials using hybrid-DFT and ML-based FFs
Simulations were performed with periodic explicit solvent models containing 64 water molecules. Importantly, including Grotthuss-type proton transfer events through MLFF sampling significantly accelerated convergence of free energy estimates. Proton insertion free energies were computed without artificial constraints, yielding an ASHEP of $-4.52\pm 0.09$ V and a real proton potential of $-11.12\pm 0.09$ eV, aligning experimental and benchmark theoretical values.
@@ IV.2.3   Screening of ether-based solvate ionic liquids
Wang et al. Wang et al. (2020) developed an active learning workflow that couples NN-MLPs with DFT calculations to efficiently screen over 1,000 oligoether ligands for their potential use in forming solvate ionic liquids (SILs) with Li+, Mg2+, and Na+. Initial ligand conformers were generated using RDKit and then optimized at the BP86-D3/def2-SVP level of theory. The NN-MLPs, trained on force and energy data from these DFT calculations, were iteratively improved via uncertainty-aware sampling of MD trajectories, allowing the model to progressively improve its predictions with each generation.
@@
For Mg2+, ligands with extended alkyl spacers and favorable coordination environments were found to offer high binding affinities, though their utility is limited by high melting points. Na+ coordination showed smaller improvements due to inherently weaker ion–dipole interactions. Interestingly, the inclusion of implicit solvent effects revealed that solvation typically reduces binding energies relative to gas-phase DFT. However, the authors noted that continuum solvation models may fall short in capturing the subtle solvation environments characteristic of SILs, making it difficult to reliably assess SIL-forming capability based on implicit solvent data alone.
@@ IV.2.5   ML-aided thermodynamics via DeepPot-SE
The method was applied to a range of aqueous species, including redox-active molecules such as hydroquinones. It achieved high chemical accuracy: redox potentials within $\sim$0.2 V and pKa values within one unit of experiment. Training required fewer than 2000 configurations per system, and the MLPs showed excellent transferability across alchemical coupling parameters.
@@ IV.2.8   Transferable GNN-based implicit solvent model for organic molecules
The GNN achieves up to 18-fold effective speedup over explicit solvent simulations (4.4$\times$ faster MD and 4$\times$ improved sampling) while matching reference free-energy profiles and intramolecular hydrogen-bond statistics. Prospective simulations reproduced conformational ensembles, torsional distributions, and solvation-driven energy barriers with high fidelity, outperforming standard implicit models. This study establishes a scalable and general-purpose ML framework for efficient solvation-aware MD in water.

### v3

@@ III.9  Error metrics
Energies are typically benchmarked against quantum chemical methods (e.g., DFT, MP2, or CCSD(T)), with mean absolute deviations (MADs) reported in meV/atom or kcal/mol. MAD below 1 kcal/mol (about $k_{B}T$ at room temperature) is generally considered chemically accurate. For dynamical simulations, however, force accuracy is more critical. Root mean square deviations (RMSDs) below 0.05 eV/Å (about 0.00097 hartree/bohr) are typically sufficient for MD simulations, vibrational spectra, and solvation structures. Since each configuration contributes $3N$ force components, force-based training provides a dense supervisory signal.
Beyond direct error metrics, models are evaluated based on their ability to reproduce experimentally observable or trajectory-derived quantities, such as solvation free energies ($\Delta G_{\text{solv}}$), RDFs, hydration numbers, or diffusivities. Both energy and force accuracies contribute to the reliability of these higher-level properties, with force fidelity crucial for trajectory-based observables.
@@ III.10  Deployment in solvation modeling
A variety of open-source tools support the training and deployment of MLPs in solvation modeling, some of which are collected in Table 6. The recent editorial survey Rupp, Küçükbenli, and Csányi (2024) highlights community-wide efforts to develop software ecosystems for atomistic ML, with a specific focus on integrating MLPs into the current tooling landscape, which includes simulation engines, workflow automation, and dataset infrastructure.
@@
[…]
@@
III.11  Sampling of configurational space
Efficient and accurate sampling of configurational space is a recurring theme across the development, training, and deployment of MLPs for solvation modeling. Case studies surveyed in Sections IV.1 and IV.2 illustrate how sampling-enhancing strategies, including active learning, metadynamics, replica exchange, and umbrella sampling, improve configurational diversity and ensure reliable prediction of solvation thermodynamics and reactivity.
Figure 12: Representative active learning or learning-on-the-fly strategies for MLP-based solvation modeling: (A) Generic active learning loop (adapted from Miksch et al. (2021)), (B) ML-accelerated MD workflow for catalytic processes (adapted from Chen, Zhang, and Zhang (2023)), (C) Energy-based structure selection for Diels–Alder reactions in water/methanol (adapted from Zhang, Juraskova, and Duarte (2024)), (D) C-NNP training at solid–liquid interfaces (adapted from Schran et al. (2021)). All subfigures are adapted from the original works for scholarly synthesis and commentary. Original copyrights remain with the respective publishers.
Sampling-enhancing techniques, namely query-by-committee (QbC) Schran et al. (2021); Schran, Brezina, and Marsalek (2020), energy-based selection Zhang, Juraskova, and Duarte (2024), and extrapolation-grade filtering Chen, Zhang, and Zhang (2023), are increasingly employed in both training and validation workflows. A recent contribution by Pan et al. Pan et al. (2025) demonstrates how pairing large-scale conformer enumeration with deep learning enables accurate tautomer ratio predictions in aqueous solution. Active learning or learning-on-the-fly (LOTF) is particularly useful for dynamically expanding datasets during MD simulations Csányi et al. (2004). Methods like DP-GEN Zhang et al. (2020), FLARE Yu et al. (2021), and AL4GAP Guo et al. (2023) rely on uncertainty estimation and targeted retraining to refine force fields during simulations systematically. C-NNP models Schran et al. (2021) and DeepPot workflows Célerse et al. (2024) exemplify successful LOTF integration.
Metadynamics, umbrella sampling, and replica exchange techniques bias sampling toward rare events such as hydrogen-bond rearrangements, proton transfer, and solvation shell exchange. The differentiability and low cost of MLPs make them ideal for coupling with these strategies. path integral molecular dynamics (PIMD) simulations further enable accurate modeling of nuclear quantum effects (NQEs) Liu, Lan, and He (2022). MLPs can be integrated into umbrella sampling schemes, where configurations are sampled under harmonic restraints along collective variables (CVs) and later combined using weighted histogram analysis methods (WHAM) Kumar et al. (1992) or multistate Bennett acceptance ratio (MBAR) Bennett (1976); Shirts and Chodera (2008). When multiple system copies are simulated at varying conditions and swapped, replica exchange molecular dynamics (REMD) further improves ensemble convergence Fabregat et al. (2020); Sugita and Okamoto (1999). Accelerated MD techniques also benefit from MLP efficiency, and CV-based approaches expand configurational reach Goetz, Cruzeiro, and Roy (2020). A conceptually distinct approach is offered by Boltzmann generators Noé et al. (2019), which use invertible NNs to learn a direct mapping between a simple latent distribution (e.g., Gaussian) and the Boltzmann-weighted configuration space, showing promise for overcoming rare event bottlenecks in solvation modeling.
A recent study by Mendible-Barreto et al. Mendible-Barreto, Whitmer, and Colón (2025) offers a detailed analysis of how the choice and distribution of CVs in training datasets affect the ability of equivariant NN-MLPs to recover free energy surface (FES) s. By systematically comparing training strategies based on classical and ab initio data for systems like butane and alanine dipeptide, the authors show that even sparse or biased CV sampling can yield accurate FES predictions if high-quality reference energies are used, thereby highlighting a practical route for sampling-aware MLP training.
Efficient and diverse sampling is essential not only during training data generation and uncertainty quantification, but also in achieving accurate predictions of solvation thermodynamics, reactivity, and interfacial behavior. These strategies enable broader configurational coverage and improved convergence of key observables in chemically and dynamically complex solvation environments. PIMD with DP-MP2 Liu, Lan, and He (2022) captures NQEs in RDFs and diffusion. However, non-conservative force-only MLPs may produce unphysical energy differences, biasing TI/FEP unless trained on both energy and force to ensure consistency: Inconsistent total energy, $U$, leads to unreliable free-energy, $F$, estimates as F=UTSF=U-TS. Hence, conservative MLPs are essential for rigorous thermodynamic modeling.
Recent work by Axelrod and Gómez-Bombarelli Axelrod and Gómez-Bombarelli (2023) systematically examines the role of conformer ensembles in molecular property prediction using deep learning, emphasizing the importance of physically realistic conformer sampling, offering practical guidance for incorporating ensemble information into ML workflow. Global structure generators (e.g., ABCluster Elm (2020)) provide diverse initial states. Smith et al. Smith et al. (2018) used QbC to efficiently train ANI-1x using only 25% of the data, while Kulichenko et al. Kulichenko et al. (2024) emphasized uncertainty-guided selection. The importance of chemically meaningful and thermodynamically diverse configurations is further emphasized by the use of force diversity metrics Botu and Ramprasad (2015), electrostatic screening Behler and Csányi (2021), dimensionality reduction Friederich et al. (2021), and benchmark diagnostics Pinheiro et al. (2021). For spectroscopically sensitive solvation (e.g., hydrogen bonding), the centroid-based self-consistent-charge density-functional tight-binding method (SCC-DFTB) has also been proposed Brezina, Beck, and Marsalek (2024).
@@ IV  Case Studies: Trends and Themes
    A2: Direct prediction of solvation-dependent properties without PES construction. In these studies, MLPs are not used to model full PESs. Instead, they are trained to predict solvation-dependent properties, such as dipole moments, redox potentials, or solvation energies, directly from the molecular structure or other structural descriptors. These approaches bypass full force field construction and focus on property-specific inference.
@@ IV.1.1  Water contact layer at solid/liquid interfaces modeled by NN-MLPs
Gäding et al. Gäding et al. (2024) applied MD simulations powered by committee neural network potentials (C-NNPs) to investigate the structure of interfacial water at six different solid surfaces: graphene, MoS2, Au(111), Au(100), Pt(111), and Pt(100). The models were built on the BPNN architecture, using ACSFs as descriptors, and were trained iteratively on reference ab initio data using an active learning protocol enhanced by a QbC approach. To monitor model reliability during training, the authors evaluated the root mean squared error (RMSE) in the potential energy per atom.
@@ IV.1.3  Metadynamics of oxygen reduction at Au–water interfaces
To ensure broad coverage of chemical space, CUR matrix decomposition, which is a low-rank approximation technique to decompose a matrix $\mathbf{A}$ into a compressed matrix product $\mathbf{C}\times\mathbf{U}\times\mathbf{R}$, was employed during dataset curation, with Euclidean distances in structural space used to identify and retain the most diverse and informative samples. The resulting model achieved energy MAEs below 1 meV/atom and force MAEs between 12.6 and 16.3 meV/Å.
@@ IV.1.4  Neural network potential for protonated water clusters in helium droplets
[…]
To incorporate NQEs, the trained NNP was coupled to ab initio path integral molecular dynamics (AI-PIMD) and path integral Monte Carlo (PIMC) simulations at 1.67 K. The AI-PIMD simulations employed the PBE exchange–correlation functional with D3 dispersion corrections for up to two-body terms. Simulations with the NN-MLP accurately captured structural observables such as spatial distribution functions (SDFs) and RDFs of helium atoms surrounding the solutes, both for microsolvated states and extended liquid helium environments.
[…]
@@
[…]
@@ IV.1.5  Solvent-driven ring opening of N-enoxyphthalimide
Célerse et al. Célerse et al. (2024) developed a DeepMD MLP using the DeepPot-SE framework and data at the hybrid-DFT-level with transfer learning to investigate solvent-mediated reactivity in bulk-phase environments. Their workflow combined active learning with well-tempered metadynamics (WT-MTD) to construct fully converged FES s for the ring-opening reaction of N-enoxyphthalimide in methanol and trifluoroethanol (TFE). These simulations uncovered a diverse ensemble of transition state (TS) configurations, often stabilized by up to five explicit solvent molecules, reflecting the inherently dynamic nature of solute–solvent interactions.
@@ IV.1.6  Explicit solvent modeling of adsorption and reactivity at metal–water interfaces
Chen et al. Chen, Zhang, and Zhang (2023) demonstrated the power of machine learning-accelerated molecular dynamics (MLaMD) with active learning to model heterogeneous catalysis under explicit solvation conditions. They employed an MTP-based linear-MLP, trained via Maxvol-based active learning. The resulting model retained DFT-level accuracy (MADs of 1 meV/atom and 0.04 eV/Å). MTP outperformed SchNet and ANI in predicting water energetics and RDFs, achieving better accuracies even with small training sets (800 DFT configurations) and enabling long-timescale MLaMD simulations of up to 500 ps compared to the AIMD simulations (50 ps).
@@ IV.1.8  Explicit solvent dynamics of the Menshutkin reaction via DeepPot-SE
WT-MetaD simulations using the trained NN-MLP allowed reconstruction of converged FES, yielding a barrier of 29.7 kcal/mol and a reaction free energy of $39\pm 2$ kcal/mol, consistent with experimental trends and outperforming SMD-based implicit solvation. FES analysis revealed significant changes in solvation structure along the reaction coordinate, particularly enhanced stabilization of the charged products via hydrogen bonding. The study highlights how MLPs trained with explicit water environments can capture both the energetic and dynamic solvation effects that are missing in continuum models.
@@
Spatial distribution functions revealed two prominent solvation caps where $\pi$-hydrogen bonds form, with comparable structures in both solvents. A time-resolved existence criterion was used to determine bond populations and lifetimes (1.7–1.8 ps). Vibrational spectroscopy (via TRPMD) captured blue-shifted OH stretching features in water, consistent with experimental Raman SC spectra. Importantly, water showed significant $\pi$-hydrogen bond anticooperativity between the two binding sites on benzene, while ammonia did not, due to weaker solvent–solvent H-bonding. Despite accurate reproduction of most properties, the C-NNP model underestimated the anticooperativity effect, highlighting limitations of local descriptors and standard training protocols. The authors suggest using MP-NN-MLPs architectures such as NequIP with customized active learning to address such long-range effects.
@@ IV.1.10  Water at solid–liquid and confined interfaces
The C-NNPs achieved high accuracy with compact training sets ($\sim$300 configurations) and were validated using an automated protocol analyzing RDFs, VDOS, and force RMSEs. For the water–TiO2 interface, long (5 ns) simulations enabled insight into interfacial layering, diffusion anisotropy, and free energy profiles, revealing firmly structured and immobile water contact layers. The framework leverages symmetry functions and fixed hyperparameters, requiring minimal user tuning, and is implemented in the open-source packages. This study exemplifies a scalable workflow for solvation-aware MLP development and application.
@@ IV.2.2  ML-aided computation of absolute redox potentials using hybrid-DFT and ML-based FFs
Simulations were performed with periodic explicit solvent models containing 64 water molecules. Importantly, including Grotthuss-type proton transfer events through MLFF sampling significantly accelerated convergence of free energy estimates. Proton insertion free energies were computed without artificial constraints, yielding an ASHEP of $4.52\pm 0.09$ V and a real proton potential of $11.12\pm 0.09$ eV, aligning experimental and benchmark theoretical values.
@@ IV.2.3  Screening of ether-based solvate ionic liquids
Wang et al. Wang et al. (2020) developed an active learning workflow that couples NN-MLPs with DFT calculations to efficiently screen over 1,000 oligoether ligands for their potential use in forming solvate ionic liquids (SILs) with Li+, Mg2+, and Na+. Initial ligand conformers were generated using RDKit and then optimized at the BP86-D3/def2-SVP level of theory. The NN-MLPs, trained on force and energy data from these DFT calculations, were iteratively improved via uncertainty-aware sampling of MD trajectories, allowing the model to improve its predictions with each generation progressively.
@@
For Mg2+, ligands with extended alkyl spacers and favorable coordination environments were found to offer high binding affinities, though high melting points limit their utility. Na+ coordination showed smaller improvements due to inherently weaker ion–dipole interactions. Interestingly, the inclusion of implicit solvent effects revealed that solvation typically reduces binding energies relative to gas-phase DFT. However, the authors noted that continuum solvation models may fall short in capturing the subtle solvation environments characteristic of SILs, making it difficult to assess SIL-forming capability based on implicit solvent data alone reliably.
@@ IV.2.5  ML-aided thermodynamics via DeepPot-SE
[…]
@@ IV.2.8  Transferable GNN-based implicit solvent model for organic molecules
[…]